OpenStudy (anonymous):

A baseball player hits a ball toward the outfield. The height h of the ball in feet is modeled by h(t) = -16t2 + 22t + 3, where t is the time in seconds. In addition, the function d(t) = 85t models the horizontal distance d traveled by the ball. If no one catches the ball, how long will it stay in the air?

OpenStudy (anonymous):

-16t2 + 22t + 3 = 0 i think

OpenStudy (anonymous):

1.5

OpenStudy (anonymous):

how did you get 1.5?

OpenStudy (anonymous):

just solve the equation

OpenStudy (anonymous):

I did, and i didn't 1.5

OpenStudy (anonymous):

what u have

OpenStudy (anonymous):

yes the roots are 1.5 and a negative root

OpenStudy (anonymous):

im also suppose to find out what the horizontal distance that the ball travels before it hits the ground? I got 127.5 can anyone tell me if im wrong?

OpenStudy (anonymous):

yes but in this problem you don't need this

OpenStudy (anonymous):

For this question, only the first equation is required

OpenStudy (anonymous):

I know, but it was the problem under it. I was suppose to find it out also. But thank you (:

OpenStudy (anonymous):

So you get the time = 1.5 d(1.5) = 85 * 1.5 = ?